# Multivariate spatial models for small area estimation of species-specific forest inventory parameters Jeffrey W. Doser1, Malcolm S. Itter2, Grant M. Domke3, Andrew O. Finley4, 5

### v1

@@ Abstract
National Forest Inventories (NFIs) provide statistically reliable information on forest resources at national and other large spatial scales. As forest management and conservation needs become increasingly complex, NFIs are being called upon to provide forest parameter estimates at spatial scales smaller than current design-based estimation procedures can provide. This is particularly true when estimates are desired by species or species groups, which is often required to inform wildlife habitat management, sustainable forestry certifications, or timber product assessments. Here we propose a multivariate spatial model for small area estimation of species-specific forest inventory parameters. The hierarchical Bayesian modeling framework accounts for key complexities in species-specific forest inventory data, such as zero-inflation, correlations among species, and residual spatial autocorrelation. Importantly, by fitting the model directly to the individual plot-level data, the framework enables estimates of species-level forest parameters, with associated uncertainty, across any user-defined small area of interest. A simulation study revealed minimal bias and higher accuracy of the proposed model-based approach compared to the design-based estimator and a non-parametric k-nearest neighbor (kNN) estimator. We applied the model to estimate species-specific county-level aboveground biomass for the 20 most abundant tree species in the southern United States using Forest Inventory and Analysis (FIA) data. Biomass estimates from the proposed model had high correlations with design-based estimates and kNN estimates. Importantly, the proposed model provided large gains in precision across all 20 species. On average across species, 91.5% of county-level biomass estimates had higher precision compared to the design-based estimates. The proposed framework improves the ability of NFI data users to generate species-level forest parameter estimates with reasonable precision at management-relevant spatial scales.
@@ 1 Introduction
Here we develop a multivariate spatial model for SAE of species-specific forest inventory parameters. Our hierarchical Bayesian model simultaneously accounts for zero-inflation in individual species responses, correlations among species, and residual spatial autocorrelation—factors that notoriously complicate analysis of forest inventory data (e.g., Finley et al., 2012; Clark et al., 2014; May et al., 2023). We implement the model at the unit-level (i.e., FIA plot-level), which provides flexibility in generating species-specific small area estimates across user-defined areal units of interest. We perform a simulation study to assess bias and accuracy of the model relative to alternative design-based and non-parametric estimators. The proposed model is applied to FIA data in the southern US to estimate county-level aboveground live and standing dead tree biomass for the 20 most abundant tree species in the region. The multivariate model generates county-level estimates with substantially improved precision compared to alternative design-based estimators, showcasing the utility of our approach to generate species-level forest parameter estimates from NFI data at management-relevant spatial scales.
@@ 2.1 FIA NFI data
On forested FIA plots, trees with a diameter at breast height (d.b.h) $\geq 12.7$ cm are measured on a cluster of four circular 0.02 ha subplots centered on each plot location (Bechtold and Patterson,, 2005). Within each subplot, trees 2.54-12.70 cm d.b.h. are measured on a circular 13.5 m2 microplot. Tree attributes (e.g., species, alive/dead/harvested status) and quantitative variables (e.g., d.b.h., height, volume) are measured following nationally-standardized protocols. Aboveground biomass for each measured tree is estimated following FIA’s National Scale Volume and Biomass Estimators (NSVB) framework (Westfall et al.,, 2024; Johnson et al.,, 2024). These individual tree estimates are summarized and expanded to provide a plot-level estimate of aboveground biomass (Mg/ha).
@@ 2.2 FIA sampling design and design-based estimators
As is commonly done in forest inventory, FIA uses post-stratified design-based estimators to estimate forest parameters of interest in an attempt to reduce variance of direct estimates without post-stratification (Westfall et al.,, 2011). Post-stratification proceeds by using wall-to-wall digital imagery to classify locations into a set of strata (e.g., forest/nonforest, canopy cover proportion classes). Stratum weights are then calculated based on the proportion of all pixels in the population assigned to the given stratum, and population estimates are obtained as a weighted average of the stratum-specific estimates (see Westfall et al., 2022 for a full description of FIA post-stratification techniques). While useful for estimating national and regional level forest parameters, such post-stratified estimators are limited in SAE applications, particularly when generating estimates by species due to individual stratum not being present within each SAE, highly zero-inflated and right-skewed distributions of species-specific forest parameters, and very small sample sizes within each stratum/SAE combination. Accordingly, for comparison with our model-based estimates, we use a simple design-based direct estimate (i.e., the sample mean). More specifically, letting $y_{i,j,k}$ denote the aboveground biomass for species $j$ in plot $i$ that falls in county $k$, our design-based estimate for county-level biomass is
@@ 2.3 Modeling framework
We seek a modeling framework that can provide improved precision of species-specific forest parameter estimates by incorporating plot-level predictor variables correlated with aboveground biomass, exploiting correlations between co-occurring species, and borrowing information from spatially proximate plots. We accomplish this via a multivariate log-normal hurdle model that accommodates the large zero-inflation and heavily right-skewed distribution of species-level biomass. The proposed model first jointly estimates the presence or absence of each species across the set of $n=46,710$ FIA plots to establish the distribution of each species across the region. We subsequently estimate species-specific biomass only at those plots where the individual species is present. We refer to these sub-models as Stage 1 and Stage 2, respectively. Estimates from the model are then used to predict species-specific biomass across a dense grid of cells over the region of interest, which we ultimately summarize to generate species-level estimates across the 1,306 counties in the southern US. In the subsequent model presentation, we focus on the case of estimating species-specific biomass in the southern US, but note that the model could be used to predict any species-specific forest parameter that takes positive-only values (e.g., basal area, volume, carbon).
@@
We first jointly estimate the presence or absence of all species to accommodate the large propensity of locations where a species does not occur (and thus has zero biomass). Let $\bm{s}_{i}$ denote the spatial coordinates of the $i^{\text{th}}$ forest plot for each of the $i=1,\dots,46,710$ plots. Define $z_{j}(\bm{s}_{i})$ as the presence (1) or absence (0) of species $j$ at plot $i$ for each of the $j=1,\dots,J$ species, where here $J=20$. We model $z_{j}(\bm{s}_{i})$ following
@@
where $\beta_{0,j}$ is the overall intercept for species $j$, $\beta_{1,j}$ and $\beta_{2,j}$ are linear and quadratic effects of minimum temperature (TMIN), $\beta_{3,j}$ and $\beta_{4,j}$ are linear and quadratic effects of maximum temperature (TMAX), $\beta_{5,j}$ and $\beta_{6,j}$ are linear and quadratic effects of total precipitation, and $\text{w}^{\ast}_{j}(\bm{s}_{i})$ is a spatially varying intercept for species $j$ at plot $i$ to accommodate residual spatial structure in the species distribution. We included minimum/maximum temperature and total precipitation to help establish individual species distributions given their importance in determining species ranges across the southern US. We include linear and quadratic effects to allow the relationships to peak at intermediate levels of one or more of the climate variables in relation to the climatic niche of the individual species. All climate data are 30-year (1981-2010) climate normals obtained from TerraClimate (Abatzoglou et al.,, 2018) using the R package climateR (Johnson,, 2024).
We model all regression coefficients and the non-spatial intercept for each species hierarchically (i.e., as random effects; Dorazio and Royle, 2005; Gelfand et al., 2005b ). More specifically, for each $t=0,\dots,p$, we model $\beta_{j,t}$ according to
@@
To complete the Bayesian specification of the model, we use vague normal priors with mean equal to 0 and variance set to 2.73 for all mean parameters. We use vague inverse gamma priors with both the shape and scale parameter set to 0.1 for all variance parameters. We set moderately informative uniform priors to the spatial decay parameters such that the effective spatial range (i.e., the distance at which the correlation between two sites falls to 0.05) is bounded between 50km and 2000km. This allows for sufficient flexibility to estimate fine-scale and/or broad-scale spatial autocorrelation. Finally, we assign standard normal priors to the lower triangle of the factor loadings matrix. Following Taylor-Rodriguez et al., (2019), we fix the diagonal elements of the factor loadings matrix to 1 and the upper triangle to 0 to ensure identifiability of the factor loadings and the spatial factors.
@@
$\begin{split}\mu_{j}(\bm{s}_{i})&=\alpha_{0,j}+\eta^{\ast}_{j}(\bm{s}_{i})+\alpha_{1,j}\cdot\text{TCC}_{i}+\alpha_{2,j}\cdot\text{VPD}_{i}+\\ &\alpha_{3,j}\cdot\text{PPT}_{j}+\alpha_{4,j}\cdot\text{ELEV}_{j}+\alpha_{5,j}\cdot\text{ELEV}^{2}_{j},\end{split}$ (9)
where $\alpha_{0,j}$ is the overall intercept for species $j$, $\alpha_{1,j}$ is the linear effect of percent tree canopy cover (TCC), $\alpha_{2,j}$ is the linear effect of average growing season vapor pressure deficit (VPD), $\alpha_{3,j}$ is the linear effect of total precipitation (PPT), and $\alpha_{4,j}$ and $\alpha_{5,j}$ are linear and quadratic effects of elevation (ELEV). We included vapor pressure deficit and total precipitation as covariates on expected biomass given their physiological importance for tree growth. Percent tree canopy cover at a 30m resolution was hypothesized to have a positive relationship with biomass, particularly with common species, and was obtained from the National Land Cover Database (NLCD; Houseman et al., 2023). Elevation was included given its correlation with distinct forest types across the southern US data. We extracted elevation data from Amazon Terrain Tiles using the elevatr R package (Hollister et al.,, 2023). The species-specific spatially varying intercept ($\eta^{\ast}_{j}(\bm{s}_{i})$) is modeled using the spatial factor NNGP approach described in (6) and (7). As in Stage 1, we fit the model with five spatial factors.
@@ 2.5 Prediction and small area estimation
2.6 Comparison to design-based and alternative non-parametric estimates
We compare county-level biomass estimates from the proposed multivariate spatial model to the design-based direct estimates (i.e., sample mean) previously defined in (1). Ideally, the proposed model will provide small area estimates that show high correlation with the design-based direct estimates, but are more precise. To assess improvements in precision, we compare the design-based coefficient of variation (i.e., the standard error divided by the sample mean) to the model-based coefficient of variation (i.e., posterior standard deviation divided by the posterior median) for biomass density estimates for each species in each county. Note that the design-based estimator and the Bayesian, model-based approach are based on two fundamentally different statistical paradigms and thus this comparison is only a rough characterization of the amount of uncertainty for the population estimate (i.e., the true biomass for each species in each county). See Gregoire, (1998) for a more complete discussion regarding the comparison of model-based and design-based estimates. As an assessment of correspondence between the design-based and model-based estimates, we calculate simple correlation coefficients for each species.
Finally, we compare the proposed model estimates with small area estimates obtained from a k-nearest neighbors (kNN) approach. The kNN approach has seen substantial traction in forest inventory mapping and estimation in large part due to its intuitive nature and relative simplicity (McRoberts et al.,, 2007). Furthermore, kNN is a multivariate approach that can simultaneously estimate species-level responses without making parametric assumptions about the underlying distribution of these responses (McRoberts,, 2012), thus making it an attractive alternative for species-level SAE. Here we use kNN to estimate county-level biomass for the 20 species of interest based on a Mahalanobis distance metric and 10 nearest neighbors. We implemented the kNN models using the yaImpute package (Crookston and Finley,, 2008). We compare our proposed model estimates to the kNN-based estimates using simple correlations.
@@ 2.7 Simulation study
We performed a simulation study to quantify the bias and accuracy of the proposed model-based approach to species-specific SAE. We simulated a single population of $41,501$ population units distributed across a $1.75\times 1.75$ km grid over the state of North Carolina. Each population unit represents a potential sampling unit. We simulated biomass of ten species across all population units to have similar characteristics to the observed FIA data (e.g., highly zero-inflated and right-skewed). From this population, we sampled 100 replicate data sets each comprising $1000$ population units selected at random from the population. For each of the replicate data sets, we calculated species-specific average biomass (Mg/ha) in 100 small areas (i.e., counties) across the state using the proposed model, design-based estimator, and the non-parametric kNN estimator. For each of the three estimation approaches, we calculated bias as the difference between the estimate and true population value for each species and county, as well as the root mean square error (RMSE) as a measure of average accuracy. Complete details on the simulation study are provided in Supplemental Information S2.

### v2

@@ Abstract
National Forest Inventories (NFIs) provide statistically reliable information on forest resources at national and other large spatial scales. As forest management and conservation needs become increasingly complex, NFIs are being called upon to provide forest parameter estimates at spatial scales smaller than current design-based estimation procedures can provide. This is particularly true when estimates are desired by species or species groups, which is often required to inform wildlife habitat management, sustainable forestry certifications, or timber product assessments. Here we propose a multivariate spatial model for small area estimation of species-specific forest inventory parameters. The hierarchical Bayesian modeling framework accounts for key complexities in species-specific forest inventory data, such as zero-inflation, correlations among species, and residual spatial autocorrelation. Importantly, by fitting the model directly to the individual plot-level data, the framework enables estimates of species-level forest parameters, with associated uncertainty, across any user-defined small area of interest. A simulation study revealed minimal bias and higher accuracy of the proposed model-based approach compared to the design-based estimator. We applied the model to estimate species-specific county-level aboveground biomass for the 20 most abundant tree species in the southern United States using Forest Inventory and Analysis (FIA) data. County-level biomass estimates from the proposed model had high correlations with design-based estimates, yet the model-based estimates tended to have a slight positive bias relative to design-based estimates, particularly for abundant and managed species. Importantly, the proposed model provided large gains in precision across all 20 species. On average across species, 91.5% of county-level biomass estimates had higher precision compared to the design-based estimates. Future work should explore incorporation of additional auxiliary data sources that can help explain fine-scale variation in biomass of managed species. The proposed framework is an attractive solution for NFI data users to generate species-level forest parameter estimates with reasonable precision at management-relevant spatial scales.
Keywords: Forest Inventory and Analysis; National Forest Inventory; Bayesian; model-based inference.
@@ 1 Introduction
Here we develop a multivariate spatial model for SAE of species-specific forest inventory parameters. Our hierarchical Bayesian model simultaneously accounts for zero-inflation in individual species responses, correlations among species, and residual spatial autocorrelation—factors that notoriously complicate analysis of forest inventory data (e.g., Finley et al., 2012; Clark et al., 2014; May et al., 2023). We implement the model at the unit-level (i.e., FIA plot-level), which provides flexibility in generating species-specific small area estimates across user-defined areal units of interest. We perform a simulation study to assess bias and accuracy of the model relative to alternative design-based estimators. The proposed model is applied to FIA data in the southern US to estimate county-level aboveground live and standing dead tree biomass for the 20 most abundant tree species in the region. The multivariate model generates county-level estimates with substantially improved precision compared to alternative design-based estimators, showcasing the utility of our approach to generate species-level forest parameter estimates from NFI data at management-relevant spatial scales.
@@ 2.1 FIA NFI data
On forested FIA plots, trees with a diameter at breast height (d.b.h) $\geq 12.7$ cm are measured on a cluster of four circular 0.02 ha subplots centered on each plot location (Bechtold and Patterson,, 2005). Within each subplot, trees 2.54-12.70 cm d.b.h. are measured on a circular 13.5 m2 microplot. Tree attributes (e.g., species, alive/dead/harvested status) and quantitative variables (e.g., d.b.h., height, volume) are measured following nationally-standardized protocols. Aboveground biomass for each measured tree is estimated following FIA’s National Scale Volume and Biomass Estimators (NSVB) framework (Westfall et al.,, 2024; Johnson et al.,, 2025). These individual tree estimates are summarized and expanded to provide a plot-level estimate of aboveground biomass (Mg/ha).
@@ 2.2 FIA sampling design and design-based estimators
As is commonly done in forest inventory, FIA uses post-stratified design-based estimators to estimate forest parameters of interest in an attempt to reduce variance of direct estimates without post-stratification (Westfall et al.,, 2011). Post-stratification proceeds by using wall-to-wall digital imagery to classify locations into a set of strata (e.g., forest/nonforest, canopy cover proportion classes). Stratum weights are then calculated based on the proportion of all pixels in the population assigned to the given stratum, and population estimates are obtained as a weighted average of the stratum-specific estimates (see Westfall et al., 2022 for a full description of FIA post-stratification techniques). While useful for estimating national and regional level forest parameters, such post-stratified estimators are limited in SAE applications, particularly when generating estimates by species due to individual stratum not being present within each SAE, highly zero-inflated and right-skewed distributions of species-specific forest parameters, and very small sample sizes within each stratum/SAE combination. Accordingly, for comparison with our model-based estimates, we use a simple design-based direct estimator (i.e., the sample mean). More specifically, letting $y_{i,j,k}$ denote the aboveground biomass for species $j$ in plot $i$ that falls in county $k$, our design-based estimate for county-level biomass is
@@
2.3 Auxiliary data
We extracted climate, forest cover, and remotely-sensed elevation data for use as auxiliary information in the subsequently described SAE framework. We extracted all climate variables from TerraClimate (Abatzoglou et al.,, 2018), which provides high-resolution ($1/24$ degree or approximately $4$km) climate normals from 1981-2010 across the globe. In particular, we extracted monthly climate normals for maximum temperature, minimum temperature, total precipitation, and vapor pressure deficit. From these monthly values, we calculated maximum temperature (TMAX) as the maximum of the monthly maximum temperature normals, minimum temperature (TMIN) as the minimum of the monthly minimum temperature normals, total precipitation (PPT) as the sum of the monthly precipitation normals, and vapor pressure deficit (VPD) as the average of the growing season (April through October) monthly vapor pressure deficit normals. We calculated tree canopy cover (TCC) at each FIA plot center in 2019 using the National Land Cover Database (NLCD) 2021 Forest Service Science data set (Houseman et al.,, 2023). Finally, we determined the elevation at the center of each FIA plot using bare-earth terrain height data from Amazon Web Service Terrain Tiles (accessed on 13 June 2025 from https://registry.opendata.aws/terrain-tiles).
@@ 2.4 Modeling framework
We seek a modeling framework that can provide improved precision of species-specific forest parameter estimates by incorporating plot-level predictor variables correlated with aboveground biomass, exploiting correlations between co-occurring species, and borrowing information from spatially proximate plots. We accomplish this via a multivariate log-normal hurdle model that accommodates the large zero-inflation and heavily right-skewed distribution of species-level biomass. The proposed model first jointly estimates the presence or absence of each species across the set of $n=46,710$ FIA plots to establish the distribution of each species across the region. We subsequently estimate species-specific biomass only at those plots where the individual species is present. We refer to these sub-models as Stage 1 and Stage 2, respectively. Estimates from the model are then used to predict species-specific biomass across a dense grid of cells over the region of interest, which we ultimately summarize to generate species-level estimates across the 1,306 counties in the southern US. In the subsequent model presentation, we focus on the case of estimating species-specific biomass in the southern US, but note that the model could be used to predict any species-specific forest parameter that takes positive-only values (e.g., basal area, volume, carbon). We used the following R packages for data downloading, processing, and manipulating auxiliary data: sf (Pebesma and Bivand,, 2023), tidyverse (Wickham et al.,, 2019), climateR (Johnson,, 2024), and elevatr (Hollister et al.,, 2023).
@@
We first jointly estimate the presence or absence of all species to accommodate the large propensity of locations where a species does not occur (and thus has zero biomass). Let $\bm{s}_{i}$ denote the spatial coordinates of the $i^{\text{th}}$ forest plot for each of the $i=1,\dots,46,710$ plots (e.g., the easting and northing coordinate for the center subplot of a FIA inventory plot $i$). Define $z_{j}(\bm{s}_{i})$ as the presence (1) or absence (0) of species $j$ at plot $i$ for each of the $j=1,\dots,J$ species, where here $J=20$. We model $z_{j}(\bm{s}_{i})$ following
@@
where $\beta_{0,j}$ is the overall intercept for species $j$, $\beta_{1,j}$ and $\beta_{2,j}$ are linear and quadratic effects of minimum temperature (TMIN), $\beta_{3,j}$ and $\beta_{4,j}$ are linear and quadratic effects of maximum temperature (TMAX), $\beta_{5,j}$ and $\beta_{6,j}$ are linear and quadratic effects of total precipitation, and $\text{w}^{\ast}_{j}(\bm{s}_{i})$ is a spatially varying intercept for species $j$ at plot $i$ to accommodate residual spatial structure in the species distribution. We included minimum/maximum temperature and total precipitation to help establish individual species distributions given their importance in determining species ranges across the southern US. We include linear and quadratic effects to allow the relationships to peak at intermediate levels of one or more of the climate variables in relation to the climatic niche of the individual species.
We model all regression coefficients and the non-spatial intercept for each species hierarchically (i.e., as random effects; Dorazio and Royle, 2005; Gelfand et al., 2005). More specifically, for each $t=0,\dots,p$, we model $\beta_{j,t}$ according to
@@
To complete the Bayesian specification of the model, we use normal priors with mean equal to 0 and variance set to 2.72 for all mean parameters, which results in a vague prior on the probability scale (Northrup and Gerber,, 2018). We use vague inverse gamma priors with both the shape and scale parameter set to 0.1 for all variance parameters. We set moderately informative uniform priors to the spatial decay parameters such that the effective spatial range (i.e., the distance at which the correlation between two sites falls to 0.05) is bounded between 50km and 2000km. This allows for sufficient flexibility to estimate fine-scale and/or broad-scale spatial autocorrelation. Finally, we assign standard normal priors to the lower triangle of the factor loadings matrix. Following Taylor-Rodriguez et al., (2019), we fix the diagonal elements of the factor loadings matrix to 1 and the upper triangle to 0 to ensure identifiability of the factor loadings and the spatial factors.
@@
$\begin{split}\mu_{j}(\bm{s}_{i})&=\alpha_{0,j}+\eta^{\ast}_{j}(\bm{s}_{i})+\alpha_{1,j}\cdot\text{TCC}_{i}+\alpha_{2,j}\cdot\text{VPD}_{i}+\\ &\alpha_{3,j}\cdot\text{PPT}_{i}+\alpha_{4,j}\cdot\text{ELEV}_{i}+\alpha_{5,j}\cdot\text{ELEV}^{2}_{i},\end{split}$ (9)
where $\alpha_{0,j}$ is the overall intercept for species $j$, $\alpha_{1,j}$ is the linear effect of percent tree canopy cover (TCC), $\alpha_{2,j}$ is the linear effect of average growing season vapor pressure deficit (VPD), $\alpha_{3,j}$ is the linear effect of total precipitation (PPT), and $\alpha_{4,j}$ and $\alpha_{5,j}$ are linear and quadratic effects of elevation (ELEV). We included vapor pressure deficit and total precipitation as covariates on expected biomass given their physiological importance for tree growth. Percent tree canopy cover was hypothesized to have a positive relationship with biomass, particularly with common species. Elevation was included given its correlation with distinct forest types across the southern US data. The species-specific spatially varying intercept ($\eta^{\ast}_{j}(\bm{s}_{i})$) is modeled using the spatial factor NNGP approach described in (6) and (7). As in Stage 1, we fit the model with five spatial factors.
@@ 2.6 Prediction and small area estimation
2.7 Model assessment and comparison to design-based estimates
We compared county-level biomass estimates from the proposed multivariate spatial model to the design-based direct estimates (i.e., sample mean) previously defined in (1). Ideally, the proposed model will provide small area estimates that show high correlation with the design-based direct estimates, but are more precise. To assess improvements in precision, we compared the design-based coefficient of variation (i.e., the standard error divided by the sample mean) to the model-based coefficient of variation (i.e., posterior standard deviation divided by the posterior median) for biomass density estimates for each species in each county. The design-based estimator and the Bayesian model-based approach rely on fundamentally different statistical paradigms, so this comparison reflects how each quantifies uncertainty for its estimate of the population parameter—in this case, species-specific biomass in each county. See Gregoire, (1998) for a more complete discussion regarding the comparison of model-based and design-based estimates. As an assessment of correspondence between the design-based and model-based estimates, we calculated simple correlation coefficients for each species. We additionally calculated the average bias between model-based county-level estimates and the design-based county-level estimates. Note that these bias estimates do not necessarily represent bias from the true, unknown species-specific county-level biomass values and rather reflect differences of the model-based estimate from the design-based estimates, which themselves could be biased.
We used four-fold cross-validation to provide an assessment of the proposed model’s ability to predict species-level small area estimates in small areas that were not used to fit the model. We randomly split the 1,306 counties into four approximately equal sets. We subsequently fit the proposed model four times, where each model was fit using the inventory plots found within three of the four sets of counties. We subsequently predicted species-specific biomass density within the counties that were not used when fitting the model and estimated bias between the resulting model-based estimates and the design-based estimates. We averaged the estimated bias value for each species across the four hold-out sets for a single measure of bias for each individual species. Our approach to generating hold out sets based on counties aligned with our desire to produce county-level estimates, and thus provided a more realistic assessment of model predictive performance compared to randomly holding out individual inventory plots.
@@ 2.8 Simulation study
In addition to model assessment in the FIA case study, we performed a simulation study to quantify the bias and accuracy of the proposed model-based approach to species-specific SAE. We simulated a single population of $41,501$ population units distributed across a $1.75\times 1.75$ km grid over the state of North Carolina. Each population unit represents a potential sampling unit. We simulated biomass of ten species across all population units following the multivariate spatial model described in Section 2.4, where we first simulated the presence/absence of each species at each population unit conditional on the climate variables and four spatial factors, then subsequently simulated the biomass for each location where the species occurs conditional on climate, elevation, tree canopy cover, and four spatial factors. The spatial factors were simulated to represent a mix of fine-scale and broad-scale spatial autocorrelation, a frequent phenomenon in forest inventory data (e.g., May and Finley,, 2025). This simulation approach resulted in data with similar characteristics to the observed FIA data (e.g., highly zero-inflated and right-skewed), with species occurring on average at approximately $\sim 29\%$ (minimum $2\%$ and maximum $87\%$) of the population units in the region.
From this population, we sampled 100 replicate data sets each comprising $1000$ population units selected at random from the population. For each of the replicate data sets, we calculated species-specific average biomass (Mg/ha) in 100 small areas (i.e., counties) across the state using the proposed model and design-based estimator. For both estimation approaches, we calculated bias as the difference between the estimate and true population value for each species and county, as well as the root mean square error (RMSE) as a measure of average accuracy. Complete details on the simulation study are provided in Supplemental Information S2.